OpenStudy (anonymous):

Giving medals to someone who can help. Complete by applying polynomial identities to complex numbers. (X+(3+5i))^2

OpenStudy (freckles):

you are suppose to multiply correct?

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

\[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

I am not sure if that is the formula.. only because it says nothing about it in my lesson

OpenStudy (freckles):

\[(a+b)^2=(a+b)(a+b) =aa+ab+ab+bb \\ =a^2+ab+ab+b^2 \\ =a^2+2ab+b^2\]

OpenStudy (anonymous):

I am just not sure if that is the formula I should use because it is not in my lesson

OpenStudy (anonymous):

And my teacher wants to see my work

OpenStudy (freckles):

what are we allowed to use then? I thought we were suppose to multiply?

OpenStudy (freckles):

honestly I don't know what complete means

OpenStudy (freckles):

in this case

OpenStudy (anonymous):

They gave me a different formula to use. Hang on ill write it out

OpenStudy (anonymous):

\[a ^{2}-b ^{2}=(\sqrt{a ^{2}}+\sqrt{b ^{2}})(\sqrt{a ^{2}}-\sqrt{b ^{2}})=(a+b)(a-b)\]

OpenStudy (freckles):

I don't think that formula applies here...

OpenStudy (freckles):

\[(x+(3+5i))^2 \\=x^2+2x(3+5i)+(3+5i)^2\]

OpenStudy (freckles):

and then you could multiply the (3+5i)^2 out too

OpenStudy (anonymous):

There is two formulas to choose from do you want me to put the other one?

OpenStudy (freckles):

We don't have to use a formula to multiply

OpenStudy (freckles):

we can just multiply

OpenStudy (freckles):

 so if we had (x+(3+5i))(x+(3+5i))